OpenStudy (lakers):

Solve by completing the square: x2 - 2x - 24 = 0 A. -4, -6 B. -4, 6 C. 4, -6 D. 4, 6

OpenStudy (badhi):

You can solve this in two ways 1. find the factors for $$(x^2-2x-24)$$ Or 2. substitute each given answer to the equation and find if it satisfies

OpenStudy (anonymous):

x^2-2x+1-1-24=0 (x-1)^2-25=0 (x-1)^2=25 \[x-1=\pm \sqrt{25}\] \[x=1\pm 5\] x=6 x=-4

OpenStudy (anonymous):

B